OpenStudy (anonymous):

10x2+11x-6 A.(10x-6)(5x+1) B.(5x+2)(6x-5) C.(x-3)(10x+2) D.(5x-2)(2x+3)

OpenStudy (nathan917):

10x^2 - 11x - 6 = 0 10 x (-6) = -60 so u need two numbers one which add to give -11 and multiply to give (-60) In this case we can use -15 and 4 so we replace the middle number by these two.. or i.e. break up the middle number .. 10x^2 - 15x + 4x -6 = 5x ( 2x - 3) + 2 ( 2x - 3) = (5x + 2)(2x + 3)

OpenStudy (anonymous):

the answer is D. (5x-2) (2x+3)

OpenStudy (anonymous):